OpenStudy (anonymous):

a student bought 1 box of crayonabd 5 reams of paper for $54. she bought 5 boxes of crayon and 3 reams of paper for $50. what is the cost of each box of crayon and each ream of paper?

OpenStudy (anonymous):

let cost of 1 box of crayons be x and cost of 1 ream of paper be y. x+5y= 54 --(1) 5x+3y=50 --(2) take (1) and multiply by 5 5x+25y=270 take the above and minus (2) 5x+25y-5x-3y=270-50 22y=220 y=10 sub y=10 to (1) x+5(10) =54 x=54-50 x=4 Cost of one box of crayons is 4 dollars One ream of paper is 10dollars

OpenStudy (anonymous):

thanks you helped me alot

OpenStudy (anonymous):

Np! try to do them yourself though. it is not difficult. just have to understand how to approach the question.